OpenStudy (anonymous):

write the equation of the line that has a y-value that increases by 4 for every x that increases by 1. it also crosses the y-axis at -2 a. y=4x-2 b. y=1/4x-2 c. y=2x+1/4 d. y=2x-4

OpenStudy (cwrw238):

slope of the line= 4 to also passes through the point(-2,0) plug these values into y-y1 = m(x-x1) and simplify

OpenStudy (anonymous):

how do you plug the values in

OpenStudy (cwrw238):

m = slope = 4, (x1,y1) = (-2,0)

OpenStudy (anonymous):

simplify ?

OpenStudy (cwrw238):

example of simplifying: y = 3(x - 4) + 8 y = 3x - 12 + 8 y = 3x - 4

OpenStudy (cwrw238):

sorry i mixed up values of x and y x1= 0 , y1 = -2 so its y - (-2) = 4(x - 0)

OpenStudy (anonymous):

Write the equation of the line that has a y-value that increases by 4 for every x that increases by 1. It also crosses the y-axis at –2.

OpenStudy (cwrw238):

rearrange the equation in my last post and you'll get one of the options